OpenStudy (anonymous):

guys help me find the concavity of f(x)= x^4-4x^3+4x^2. i kow the procedure but have problems with calculations. so helpp

OpenStudy (across):

Differentiate this function twice.

OpenStudy (ash2326):

f(x)=x^4-4x^3+4x^2 f'(x)=4x^3-12x^2+8x f''(x)=12x^2-24x+8 let's find the inflection points where the curve changes concavity f''(x)=0 x=1+ 1/root 3 and 1-1/root 3 so x =0.42 and 1.577 here the curve changes concavity if f''(x)>=0 then f(x) is concave up if f''(x)<=0 then f(x) is concave down for the interval (-infinity, 0.42] and [1.577, infinity) f(x)''>=0 for the interval[0.42,1.577] f''(x)<=0 so f(x) has concavity up in the interval(-infinity, 0.42] and [1.577, infinity) and concavity down in [0.42, 1.577]

OpenStudy (anonymous):

is it correct to get intervals in decimal??

OpenStudy (ash2326):

yeah, you can also write in radical like 1+1/root 3

OpenStudy (anonymous):

plz help me with this f(x)=x^3/(3x^2+1). for the same thing. i messed up while solving.

OpenStudy (anonymous):

 First step find the first derivative